Using Gauss-Jordan elimination- solve following system of equations.
x+y+z=2
2x-3y+z=-11
-x+2y-z=8
I think it is x= 1/3, y=10/3, z=-5/3
Am I correct?
Thanks

Answer by venugopalramana(3286)

Using Gauss-Jordan elimination- solve following system of equations.
x+y+z=2
2x-3y+z=-11
-x+2y-z=8
I think it is x= 1/3, y=10/3, z=-5/3
Am I correct?
Thanks
I UNDERSTAND THAT YOU SOLVED AND GOT THE ANSWER ...THEN YOU ARE
10000% CORRECT.GOOD JOB KEEP IT UP